# Persuasion in Networks: Can the Sender Do Better than Using Public Signals?

### v2

@@ 2.3 Public signals
Again, I describe the game with $n$ receivers, and sender’s optimal payoff facing infinitely many receivers is captured as the limit of her optimal payoffs in these finite games [^8] This is equivalent as just letting the sender face a continuum of receivers, of which $\gamma_{h}$ has type $h$. I describe sender’s problem as the sequence of finite games to be consistent with the setup above for network signals.
First, the sender designs and commits to signal structure $\pi(.)$. After seeing sender’s strategy, receivers choose their action upon observing each signal and $\emptyset$. This is captured by function $a_{i}:S\cup\emptyset\rightarrow\{0,1\}$.
The nature draws a state, and then a signal is realised according to $\pi(.)$. All receivers observe the realised signal. We study the subgame perfect equilibrium of this game. We can find sender’s optimal subgame perfect equilibrium payoff for each game with $n$ receivers, and I use $V^{*}_{pub}$ to denote the limit of sender’s optimal payoffs using public signals. The main research question of this paper is to compare $V^{*}_{pub}$ and limiting network signal payoffs defined in the previous subsection.
@@ 2.4 Discussion of Setup
Empty Signal: From description of the public signal game, the empty signal’s role is not different from non-empty ones, so it is actually WLOG to assume that sender does not use empty signals in public signal game. For the network signal game, as we will see below, even if sender cannot use $\emptyset$, she can still essentially send no signal to almost everyone. Intuitively, this is done by choosing the seeds to be very isolated nodes. Allowing the sender to send $\emptyset$ directly is mainly for exposition. This also implies that one can think of $\emptyset$ as the sender choosing to conceal some signals. This interpretation requires sender to commit to concealment of signals before signal realisation. Relaxing the ability to commit to concealment plan restricts what sender can do using network signals, and thus strengthens results on public signals being better. For results on network signals being better, in the proofs for them, the sender uses strategies where non-empty signals give the sender a weakly higher payoff than $\emptyset$, so the sender has no incentive to conceal them.
3 Analysis of Receivers’ Posterior and the Network
In order to compare sender’s payoff from public and network signals, we first need to understand receivers’ best response facing sender’s strategies. This is straightforward for public signals, which is essentially the same as standard Bayesian persuasion. For network signals, analysis is more complicated since signals spread endogenously on networks. I present some lemmas that can be used to characterise limiting probabilities with which receivers observe each signal and $\emptyset$ when $n$ gets large. These lemmas can then be used to approximate and bound sender’s payoffs using network signals in large networks.
@@ 3 Analysis of Receivers’ Posterior and the Network
Given sender’s strategy $(\pi,Z)$ in game with $n$ receivers, posterior of a receiver with type $t$ and degree $d$ upon observing some signal $s\neq\emptyset$ is:
@@
The main idea of the proof is that network formation process and seeding strategy do not depend on the state, so the posterior upon observing some non-empty signal is determined completely by signal structure $\pi(.)$ and prior.
There are some important observations that follow from this lemma. If some signal $s$ persuades both types when sent publicly, then all receivers will share it if they receive it. If such a signal is realised, then a receiver will observe it if and only if she is connected to one of the seeds. In other words, all receivers on components [^9] A component in a network is a set of nodes such that all nodes in this set are connected to each other, and all nodes not in this set are not connected to nodes in the set. that contain at least one seed will observe the signal. If $s$ persuades only the believers but not the skeptics when sent publicly, then a receiver $i$ observes the signal if and only if there exists a path from $i$ to some seed such that all receivers on that path excluding $i$ are believers. In other words, all believers on components that contain at least one seed in the subnetwork of believers [^10] The subnetwork of believers is the network formed by deleting all sceptics and links from them will observe the signal, and a skeptic observes it if she has a neighbour who is a believer and observes the signal.
@@
Let $L_{i}(g^{n})$ denote the $i$-th largest component in realised network $g^{n}$. Theorem 3.1 and 3.6 in Bollobás et al. (2007) show that
@@
As argued above, if a signal is only passed on by believers, we need to look at the subnetwork of believers. I use $\hat{L}_{i}(g^{n})$ to denote the $i$-th largest component in the believers’ subnetwork of $g^{n}$. The believers’ subnetwork depends only on connections between believers. Also, conditional on connectedness of a pair of believers, whether they are connected is independent of how sceptics are connected to each other and how sceptics are connected to believers. Therefore, the distribution over connections between believers, and thus the believers’ subnetwork, will not change if we set $\lambda_{i}$ to be 0 for all sceptics. Applying 1 then gives
@@
There exists $\hat{c}\in[0,1]$ such that $|\hat{L}_{1}|/n$ converges in probability to $\hat{c}$, and $|\hat{L}_{i}|/n$ converges in probability to 0 for $i>1$.
Using these results on giant component and Lemma 1, we can specify some conditions under which the receivers observe $\emptyset$ with very high probability in large social networks.
@@
Fix any $\epsilon>0$. For $n$ large enough, given any $s\in S$ and any seeding with no seeds on $L_{1}$, the proportion of receivers observing $s$ is larger than $\epsilon$ with probability at most $\epsilon$.
Also, if $s$ is only passed on by believers and no seeds are on $\hat{L}_{1}$, the proportion of receivers observing $s$ is larger than $\epsilon$ with probability at most $\epsilon$.
@@
Fix any $\epsilon>0$. For $n$ large enough, given any $s\in S$ and any seeding, the proportion of receivers who are not on $L_{1}$ and observe $s$ is larger than $\epsilon$ with probability at most $\epsilon$.
Also, if $s$ is only passed on by believers, the proportion of receivers who have no neighbour on $\hat{L}_{1}$ and observe $s$ is larger than $\epsilon$ with probability at most $\epsilon$.
@@
Combining Lemma 3 and Lemma 2, the probability that a receiver can observe $s$ is approximately the probability that $s$ hits the appropriate giant component, times the probability that the receiver is on (or has a neighbour on) that giant component. The next lemma describes probabilities of being connected to $L_{1}$ and $\hat{L}_{1}$, conditional on type and degree of a node.
@@
There exist functions $\zeta(t,d)$ and $\hat{\zeta}(t,d)$ both increasing in $d$ such that, conditional on having degree $d$ and type $t$:
@@
    the proportion of receivers on $L_{1}$ converges in probability to $\zeta(t,d)$.
@@
    the proportion of receivers with a neighbour on $\hat{L}_{1}$ converges in probability to $\hat{\zeta}(t,d)$.
The equations which characterise $\zeta(\cdot)$ and $\hat{\zeta}(\cdot)$ are not key to the results below, so are only presented in the appendix.
@@ 4.1 Baseline network
In this subsection, I study the simplest case where the network formation process ignores the types of receivers. Formally, I study the case where $f_{t}(\cdot)$ is independent of $t$ and $q=1$.
The next lemma provides the main intuition underlying this subsection.
@@
If $f_{l}(\lambda)=f_{h}(\lambda)$ for all $d$ and $q=1$, then we have $\zeta(h,d)=\zeta(l,d)$ and $\hat{\zeta}(h,d)=\hat{\zeta}(l,d)$ for all $d$.
This follows from the fact that, in this baseline network, the network formation process essentially ignores types of receivers. Intuitively, this says that if a signal gets viral, then believers and skeptics have the same probability of receiving it when $n$ is large. Therefore, when $n$ gets large, the probability of observing a signal is not different between the two types of receivers.
@@
If $f_{l}(d)=f_{h}(d)$ for all $d$ and $q=1$, then
  1. 1.
    If $v(.)$ is linear or convex, $\limsup_{n\rightarrow\infty}V^{*}_{net}(n)\leq V^{*}_{pub}$.
  2. 2.
    If $\hat{c}>0$ (i.e. the giant component in type $h$ subnetwork is non-vanishing), there exist concave $v(.)$ such that $\liminf_{n\rightarrow\infty}V^{*}_{net}(n)>V^{*}_{pub}$.
For the second point, the following remark provides such a concave function $v(\cdot)$. It intuitively says that the sender prefers network signals if her risk aversion is high enough.
Remark 2.
Suppose that $v(x)=(x^{\beta}-1)/\beta$, which has constant relative risk aversion. If $\hat{c}>0$, the sender prefers network signals for sufficiently small $\beta$.
A necessary and sufficient condition for $\hat{c}>0$ can be found in Bollobás et al. (2007) and Akbarpour et al. (2020). A sufficient condition is that $\lambda_{min}^{2}\gamma_{h}\sum_{\lambda>0}f_{h}(\lambda)>1$, where $\lambda_{min}$ is the minimum non-zero value $\lambda$ could take.
@@
As discussed above, heterogeneity of priors introduces the trade-off between persuading skeptics and dissuading believers, and one might expect that use of network signals and possibility of seeding could alleviate this trade-off. For instance, if a signal persuades both types of receivers, ideally the sender wants believers not to observe it, and skeptics to observe it very often. Is there a way of choosing seeds to achieve this?
@@
By Lemma 2, for a signal to reach a large fraction of the population, we need the signals to be viral, i.e. the seeds of the signal needs to be on the relevant giant component. However, once the signals get viral, it is difficult to control who the signal reaches in a way that benefits the sender, unless the network exhibits certain features explored in the next subsection. In particular, in the baseline network, Lemma 5 states that once a signal gets viral, it must reach believers and sceptics with the same probability. Therefore, network signals cannot make sceptics observe informative signals more often than believers.
@@
The reason for why sender could prefer network signals for some concave payoff function is as explained in the introduction. With public signals, all receivers of the same type must be choosing the same action. With network signals, receivers of the same type can have different observations and choose different actions. Roughly speaking, public signals will persuade either all receivers or no receivers (of the same type), whereas network signals allow the sender to always persuade some receivers. A sender who is very risk averse prefers the latter. The importance of the sender’s risk attitude has been noted by Arieli and Babichenko (2019). They relate optimality of public signals to the core of cooperative games. Their results imply that if receivers have the same prior and sender’s payoff is convex, then sender’s optimal payoff can be achieved using public signals. In contrast with this paper, they also find that when the sender’s payoff is strictly increasing [^11] Their condition is actually weaker. They only need the sender to prefer 1 receiver choosing action 1 to no receiver choosing action 1, public signals cannot be optimal if receivers have heterogeneous priors, while the sender prefers public signals even with heterogeneous priors in Proposition 1. The difference is mainly caused by Arieli and Babichenko (2019) having no communication between receivers. Communication between receivers impose restrictions on correlations of observations between different receivers, which reduces the sender’s payoff using personalised signals.
@@
I provide a rough sketch of the proof for point 1. Suppose for simplicity that sender’s signal structure $\pi(\cdot)$ uses only 2 non-empty signals $s$ and $s^{\prime}$, where $s$ persuades both types of receivers to take action 1 and $s^{\prime}$ persuades only the believers, i.e. $\frac{\pi(s|\omega=1)}{\pi(s|\omega=0)}>\frac{\mu_{l}(\omega=0)}{\mu_{l}(\omega=1)}$, and $\frac{\mu_{h}(\omega=0)}{\mu_{h}(\omega=1)}<\frac{\pi(s^{\prime}|\omega=1)}{\pi(s^{\prime}|\omega=0)}<\frac{\mu_{l}(\omega=0)}{\mu_{l}(\omega=1)}$.
@@
By Lemma 2, if no seeds are on the relevant giant component, then very few people will observe it. Therefore, for the sketch here, suppose that the seeds for $s$ are always on $L_{1}$ and the seeds for $s^{\prime}$ are always on $\hat{L}_{1}$. Let $a^{*}(\emptyset;t,d,\Pi)$ denote receiver’s optimal action upon observing $\emptyset$, given receiver’s degree $d$, type $t$ and sender’s strategy $\Pi$.
@@
When $n$ gets large, roughly a proportion $\gamma_{t}p_{d}^{t}$ of the population has type $t$ and degree $d$. Among them, roughly a proportion $\zeta(t,d)$ are on the giant component $L_{1}$. Therefore, when $s$ is realised, roughly $\zeta(t,d)$ of those with type $t$ and degree $d$ observe $s$, and others choose $a^{*}(\emptyset;t,d,\Pi)$. Also, roughly a proportion $\hat{\zeta}(t,d)$ of them have a neighbour on the giant component in the subnetwork of type $h$ nodes. Therefore, when $s^{\prime}$ is realised, roughly $\hat{\zeta}(t,d)$ of those with type $t$ and degree $d$ observe $s^{\prime}$, and others choose $a^{*}(\emptyset;t,d,\Pi)$. By Lemma 5, $\zeta(t,d)$ and $\hat{\zeta}(t,d)$ don’t depend on $t$, so I omit the argument $t$ in this sketch. Also, $p_{d}^{h}$ and $p_{d}^{l}$ are the same in the baseline network so I omit the superscript. Using these, sender’s payoff is close to
@@
Therefore, to show that public signals are better, it’s sufficient to show that we can imitate the payoffs $V(d,\Pi)$ for each $d$. Consider a public signal strategy that sends the signal $s$ with probability $\pi(s|\omega)\zeta(d)$ and $s^{\prime}$ with probability $\pi(s^{\prime}|\omega)\hat{\zeta}(d)$ in each state $\omega$, and sends $\emptyset$ with the remaining probability. Note that these are roughly the probabilities that agents with degree $d$ observe $s$, $s^{\prime}$ and $\emptyset$ under the network signal strategy $\Pi$, therefore posteriors for each observation are also close to those under $\Pi$. A conjecture is that this implies that receiver’s action given each observation will be the same as those with degree $d$ under network signal strategy $\Pi$. A technical difficulty is that receiver’s action changes discontinuously at posterior 0.5, so posterior being close does not directly imply that they choose the same action. This is dealt with in the proof, and here for the sake of illustration let us take this conjecture to be true. Given this conjecture, one can check that sender’s payoff from this public signal structure is $V(d,\Pi)$.

### v3

@@ 2.3 Public signals
Again, I describe the game with $n$ receivers. The sender’s optimal payoff facing infinitely many receivers is captured as the limit of her optimal payoffs in these finite games when $n\rightarrow\infty$ [^10] This is equivalent as just letting the sender face a continuum of receivers, of which $\gamma_{h}$ has type $h$. I describe sender’s problem as the sequence of finite games to be consistent with the setup above for network signals.
First, the sender designs and commits to signal structure $\pi$. After seeing sender’s strategy, receivers choose their strategies $a_{i}:S\cup\emptyset\rightarrow\{0,1\}$, where $a_{i}(s)$ describes receiver $i$’s action upon observing $s\in S\cup\emptyset$. The nature draws a state, and then a signal is realised according to $\pi$. All receivers observe the realised signal.
In the subgame following $\pi$, receivers best respond to $\pi$ and to each other. With public signals, all receivers directly receive the same signal from the sender, so the network and $\mathbf{a}_{-i}$ are irrelevant for forming posterior. Let $\psi_{pub}(s;\pi,t_{i})$ denote posterior of receiver $i$ with type $t_{i}$. Equilibrium in subgames following sender’s strategies are similarly defined
Definition 2.
In a public signals game, receiver $i$’s strategy $a_{i}$ is a best response to $\pi$ and $\mathbf{a}_{-i}$ if $a_{i}(s)=1\iff\psi_{pub}(s;\Pi,t_{i})\geq 0.5$ for all $s\in S\cup\{\emptyset\}$. In the subgame following $\pi$, receivers’ strategy profile $\mathbf{a}$ is an equilibrium if $a_{i}$ is best responding to $\pi$ and $\mathbf{a}_{-i}$ for all $i$.
A result analogous to remark 2 holds for public signals. Given receivers’ equilibrium behaviour in subgames, the sender chooses $\pi$ to maximise expected payoff. We can find the sender’s optimal payoff for the game with $n$ receivers, denoted by $V^{*}_{pub}(n)$. The sender’s optimal payoff facing infinitely many receivers is captured by $V^{*}_{pub}\equiv lim_{n\rightarrow\infty}V^{*}_{pub}(n)$.
The main research question of this paper is to compare $V^{*}_{pub}$ and limiting payoff using network signals defined in the previous subsection.
It is worth re-emphasising that receivers’ actions are conditioned on the signal (which can be empty) they receive in the public signals game. Similarly, their actions are conditioned on the signal they receive and their degree in the network signals game. As can be seen in the definitions of equilibrium in subgames, receivers optimise their action given their signal and degree if relevant. Therefore, although receivers choose strategies before realisation of signal and network, it is equivalent to letting them move after observing their signal and degree. This arrangement of timing is standard in Bayesian persuasion and when analysing Bayesian games in general.
@@ 2.4 Discussion of Setup
Empty Signal: From description of the public signals game, the empty signal’s role is not different from non-empty ones, so it is actually WLOG to assume that sender does not use empty signals in public signals game. For the network signals game, as we will see below, even if sender cannot use $\emptyset$, she can still essentially send no signal to almost everyone. Intuitively, this is done by choosing the seeds to be very isolated nodes. Allowing the sender to send $\emptyset$ directly is mainly for exposition. This also implies that one can think of $\emptyset$ as the sender choosing to conceal some signals. This interpretation requires sender to commit to concealment of signals before signal realisation. Relaxing the ability to commit to concealment plan restricts what sender can do using network signals, and thus strengthens results on public signals being better. For the results on network signals being better, in their proofs, the sender uses strategies where non-empty signals give the sender a weakly higher payoff than $\emptyset$, so the sender has no incentive to conceal them.
3 Analysis of the Network and Receivers’ Posterior
In order to compare sender’s payoff from public and network signals, we first need to understand equilibrium in subgames given sender’s strategies. This is straightforward for public signals, which is essentially the same as standard Bayesian persuasion. For network signals, analysis is more complicated since signals spread endogenously on networks. I present some lemmas that can be used to characterise limiting probabilities with which receivers observe each signal and $\emptyset$ when $n$ gets large. These lemmas can then be used to approximate and bound sender’s payoffs using network signals in large networks.
@@ 3 Analysis of the Network and Receivers’ Posterior
In the network signals game with $n$ receivers, suppose that sender’s strategy is $\Pi=(\pi,Z)$ and other receivers’ strategy profile is $\mathbf{a}_{-i}$ in the subgame following $\Pi$. Conditional on having degree $d$ and observing $s\neq\emptyset$, receiver $i$ with type $t$ believes state $\omega=1$ realises with posterior probability
@@
The main idea of the proof is that network formation process and seeding strategy do not depend on the state, so the posterior upon observing some nonempty signal is completely determined by the signal structure $\pi$ and prior.
Some important observations follow from this lemma. If some signal $s$ persuades both types when sent publicly, then all receivers will share it if they receive it. If such a signal is realised, then a receiver will observe it if and only if she is connected to one of the seeds. In other words, all receivers on components [^11] A component in a network is a set of nodes such that all nodes in this set are connected to each other, and all nodes not in this set are not connected to nodes in the set. that contain at least one seed will observe the signal. If $s$ persuades only the believers but not the sceptics when sent publicly, then a receiver $i$ observes the signal if and only if there exists a path from $i$ to some seed such that all receivers on that path excluding $i$ are believers. In other words, if a component in the subnetwork of believers [^12] The subnetwork of believers is the network formed by deleting all sceptics and links of sceptics from $g^{n}$ contains a seed, then all believers on this component observe the signal. A sceptic observes it if she has a neighbour who is a believer and observes the signal.
@@
Let $L_{i}(g^{n})$ denote the $i$-th largest component in realised network $g^{n}$, and $|L_{i}(g^{n})|$ denote the number of nodes it contains. Theorem 3.1 and 3.6 in Bollobás et al. (2007) show that
@@
As argued above, if a signal is only passed on by believers, we need to look at the subnetwork of believers, i.e. the network formed by deleting all sceptics and their links from $g^{n}$. I use $\hat{L}_{i}(g^{n})$ to denote the $i$-th largest component in the subnetwork of believers. The subnetwork of believers depends only on links between believers. Also, conditional on connectedness of a pair of believers, whether they are linked is independent of how sceptics are linked to each other and how sceptics are linked to believers. Therefore, the distribution over links between believers, and thus the subnetwork of believers, will not change if we set $\lambda_{i}$ to be 0 for all sceptics. Applying 1 then gives
@@
There exists $\hat{c}\in[0,1]$ such that $|\hat{L}_{1}(g^{n})|/n$ converges in probability to $\hat{c}$, and $|\hat{L}_{i}(g^{n})|/n$ converges in probability to 0 for $i>1$.
Using these results on giant component and Lemma 1, we can specify some conditions under which the receivers observe $\emptyset$ with very high probability in large social networks. The following lemma shows that in large networks, if the seeds are not on the relevant giant component, then the proportion of receivers who observe the signal is very small.
@@
Fix any $\epsilon>0$. For $n$ large enough, given any $s\in S$ and any seeding strategy $Z$ such that $Z(g^{n},s)\cap L_{1}(g^{n})=\emptyset$ for all possible $g^{n}$, the proportion of receivers who observe $s$ is larger than $\epsilon$ with probability at most $\epsilon$.
Also, if $s$ is only passed on by believers and $Z(g^{n},s)\cap\hat{L}_{1}(g^{n})=\emptyset$ for all possible $g^{n}$, the proportion of receivers who observe $s$ is larger than $\epsilon$ with probability at most $\epsilon$.
@@
Fix any $\epsilon>0$. For $n$ large enough, given any $s\in S$ and any seeding strategy $Z$, the proportion of receivers who are not on $L_{1}(g^{n})$ and observe $s$ is larger than $\epsilon$ with probability at most $\epsilon$.
Also, if $s$ is only passed on by believers, the proportion of receivers who have no neighbour on $\hat{L}_{1}(g^{n})$ and observe $s$ is larger than $\epsilon$ with probability at most $\epsilon$.
@@
Combining Lemma 3 and Lemma 2, the probability that a receiver can observe $s$ is approximately the probability that $s$ hits the appropriate giant component, times the probability that the receiver is on (or has a neighbour on) that giant component. The next lemma describes probabilities of being linked to the giant component, conditional on type and degree of a node.
@@
There exist functions $\zeta(t,d)$ and $\hat{\zeta}(t,d)$ both increasing in $d$ such that, among receivers with degree $d$ and type $t$:
@@
    the proportion of those who are on $L_{1}(g^{n})$ converges in probability to $\zeta(t,d)$.
@@
    the proportion of those who have a neighbour on $\hat{L}_{1}(g^{n})$ converges in probability to $\hat{\zeta}(t,d)$.
The functions $\zeta$ and $\hat{\zeta}$ characterise limiting conditional probabilities of being on or having a neighbour on the giant components $L_{1}(g^{n})$ and $\hat{L}_{1}(g^{n})$. The equations which characterise functions $\zeta$ and $\hat{\zeta}$ are not key to the results below, so are only presented in the appendix.
Sender’s optimal strategy:
Characterisation of optimal strategies is not required for the main results below, and also does not provide much insight and intuition for them. Therefore, a brief discussion of them is relegated to Section 5 and the formal details to the online appendix.
@@ 4.1 Baseline network
In this subsection, I study the simplest case where the network formation process ignores the types of receivers. Formally, I study the case where distribution of connectedness $f_{t}$ is independent of $t$ and $q=1$. The next lemma provides part of the main intuition underlying this subsection.
@@
If $f_{l}(\lambda)=f_{h}(\lambda)$ for all $\lambda$ and $q=1$, then we have $\zeta(h,d)=\zeta(l,d)$ and $\hat{\zeta}(h,d)=\hat{\zeta}(l,d)$ for all $d$.
This follows from the fact that, in this baseline network, the network formation process essentially ignores types of receivers. Intuitively, this says that if a signal gets viral, then believers and sceptics have roughly the same probability of receiving it when $n$ is large.
@@
In networks without homophily or correlation between degree and type, a risk-neutral or risk-loving sender prefers public signals.
Formally, in networks where $f_{l}(\lambda)=f_{h}(\lambda)$ for all $\lambda$ and $q=1$, if the sender’s payoff function $v$ is linear or convex, then $\limsup_{n\rightarrow\infty}V^{*}_{net}(n)\leq V^{*}_{pub}$.
@@
As discussed in introduction, with heterogeneity of priors, the sender wants to provide information only to the sceptics but not to the believers. For instance, if a signal persuades both types of receivers, ideally the sender wants believers not to observe it, and sceptics to observe it very often. Is there a way of choosing seeds to achieve this using network signals?
@@
By Lemma 2, for a signal to reach a large fraction of the population, we need the signals to be viral, i.e. the seeds of the signal needs to be on the relevant giant component. However, once the signals become viral, it is difficult to control who the signal reaches. In particular, in the baseline network, Lemma 5 states that once a signal becomes viral, it must reach believers and sceptics with the same probability. Therefore, network signals cannot make sceptics observe informative signals more often than believers.
@@
Arieli and Babichenko (2019) finds that when receivers are homogeneous, public signals are optimal for the sender. However, in sharp contrast with Proposition 1 above, they also find that when the sender’s payoff function is strictly increasing [^13] Their condition is actually weaker. They only need the sender to prefer 1 receiver choosing action 1 to no receiver choosing action 1, public signals cannot be optimal if receivers are heterogeneous, while the sender prefers public signals even with heterogeneous priors in Proposition 1. The difference is driven by Arieli and Babichenko (2019) having no communication between receivers. Communication between receivers impose restrictions on correlations of observations between different receivers. Lemma 5 sheds light on these restrictions, and intuition above explains why these restrictions imply that network signals cannot outperform public signals.
@@
I provide a rough sketch of the proof for Proposition 1. Suppose for simplicity that the sender’s network signal structure $\pi$ uses only 2 non-empty signals $s$ and $s^{\prime}$, where $s$ persuades both types of receivers to take action 1 and $s^{\prime}$ persuades only the believers, i.e. $\frac{\pi(s|\omega=1)}{\pi(s|\omega=0)}>\frac{\mu_{l}(\omega=0)}{\mu_{l}(\omega=1)}$, and $\frac{\mu_{h}(\omega=0)}{\mu_{h}(\omega=1)}<\frac{\pi(s^{\prime}|\omega=1)}{\pi(s^{\prime}|\omega=0)}<\frac{\mu_{l}(\omega=0)}{\mu_{l}(\omega=1)}$.
@@
By Lemma 2, if no seeds are on the relevant giant component, then very few people will observe it. Therefore, for the sketch here, suppose that the seeds for $s$ are always on $L_{1}(g^{n})$ and the seeds for $s^{\prime}$ are always on $\hat{L}_{1}(g^{n})$. For receivers with degree $d$ and type $t$, let $a^{*}(\emptyset;t,d,\Pi)$ denote their action upon observing $\emptyset$ in equilibrium of the subgame following sender’s strategy $\Pi$.
@@
When $n$ gets large, roughly a proportion $\gamma_{t}p_{d}^{t}$ of the population has type $t$ and degree $d$. Among them, roughly a proportion $\zeta(t,d)$ is in the giant component $L_{1}$. Therefore, when $s$ is realised, roughly $\zeta(t,d)$ of those with type $t$ and degree $d$ observe $s$, and others choose $a^{*}(\emptyset;t,d,\Pi)$. Also, roughly a proportion $\hat{\zeta}(t,d)$ of them have a neighbour on the giant component in the subnetwork of type $h$ nodes. Therefore, when $s^{\prime}$ is realised, roughly $\hat{\zeta}(t,d)$ of those with type $t$ and degree $d$ observe $s^{\prime}$, and others choose $a^{*}(\emptyset;t,d,\Pi)$. By Lemma 5, $\zeta(t,d)$ and $\hat{\zeta}(t,d)$ do not depend on $t$, so I omit the argument $t$ in this sketch. Also, $p_{d}^{h}$ and $p_{d}^{l}$ are the same in the baseline network, so I omit the superscript. Using these, sender’s payoff from $\Pi$ is close to
@@
Therefore, to show that public signals are better, it’s sufficient to show that we can imitate the payoffs $V(d,\Pi)$ for each $d$. Consider a public signal strategy that sends the signal $s$ with probability $\pi(s|\omega)\zeta(d)$ and $s^{\prime}$ with probability $\pi(s^{\prime}|\omega)\hat{\zeta}(d)$ in each state $\omega$, and sends $\emptyset$ with the remaining probability. Note that these are roughly the probabilities that agents with degree $d$ observe $s$, $s^{\prime}$ and $\emptyset$ under the network signal strategy $\Pi$, therefore posteriors for each observation are also close to those under $\Pi$. A conjecture is that this implies that receiver’s action given each observation will be the same as those with degree $d$ under the network signal strategy $\Pi$. A technical difficulty is that receiver’s action changes discontinuously at posterior 0.5, so posterior being close does not directly imply that they choose the same action. This is dealt with in the proof, and here for the sake of illustration, let us take this conjecture to be true. Given this conjecture, one can check that the sender’s payoff from this public signal structure is $V(d,\Pi)$.